OpenStudy (anonymous):

If y is directly proportional to x^2 and y =9 when x =2, what is y when x=4

OpenStudy (solomonzelman):

" y is directly proportional to x^2" y=k x^2 'y =9 when x =2" 9= k (2)^2 9=4k k=9/4 what is y when x=4?

OpenStudy (anonymous):

would I multiply 9*4? 36?

OpenStudy (anonymous):

basically, after simplifying that 9/4 * 4^2 is same as 9*4. Correct.